Hi! Can someone explain how to to answer the problem? I have tried but I am confused on how to write it?


Assume that a competitive market is currently represented by the following supply and demand equations: Qs P 20 Qd 110-0.5P 1. (2 Points) Solve for the equilibrium price and quantity. 2. (1 Point) Solve for the price elasticity of demand. 3. (4 Points) Graph the market equilibrium. Fully label the graph.
Answer #1

1) For equilibrium in the market we have

Qd = Qs

110 - 0.5P = P - 20

110 + 20 = 0.5P + P

P = P* = 130/1.15 = 86.67

Qs = Qd = Q* = 66.67

2) Price elasticity of demand = ed = slope of demand function x P*/Q* = -0.5 x 86.67 / 66.67 = -0.65

3) Graph is shown below

240 210 180 150 2 120 90 86.67 60 30 0 66.67 0 20 40 60 80100 120 140160 180 200 Quantity Demand Supply